OpenStudy (anonymous):

Suppose two sides of an equilateral triangle are represented by 5y - 6 and 4y + 12. What is the value of y?

OpenStudy (anonymous):

Equilateral means the sides are the same size. That means you can put those two equations equal to each other. 4y + 12 = 5y - 6 18 = y

OpenStudy (anonymous):

5y-6=4y+12

OpenStudy (anonymous):

So whats the value of Y?

OpenStudy (anonymous):

y= 18